OpenStudy (anonymous):

what is 3^n+4 = 27^2n

OpenStudy (anonymous):

im trying to solve for x by determining the common base

OpenStudy (anonymous):

then itd be (3^n)+4=3^6n

OpenStudy (anonymous):

27 is 3^3 so 27^2n is 3^6n then its just n+4=6n

OpenStudy (anonymous):

n=4/5

OpenStudy (anonymous):

wait is it 3^(n+4) or 3^n + 4

OpenStudy (anonymous):

i want to find the common base,

OpenStudy (anonymous):

I assumed it was 3^n+4

OpenStudy (anonymous):

ya thats just a much harder problem depends on what math class this is for

OpenStudy (anonymous):

if the problem is like 3^(n+4) then it can be easily solved if its like 3^n+4 then it'd be done by log properties.

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

im looking for n

OpenStudy (anonymous):

could you please correctly state the problem using paranthesis if necessary.

OpenStudy (anonymous):

its either 4/5 or .8ln(1/3)

OpenStudy (anonymous):

depending on if the problem does or does not have parenthesis respectively

OpenStudy (anonymous):

solve for x by determind a common base showing all your work : 3^n+4 = 27^2n

OpenStudy (anonymous):

if you would like help could you please attach the paper that has the problem on it for us to analyze what is the correct way the problem is written?

OpenStudy (anonymous):

is the problem: \[3^{n+4} \] or \[3^n+4\]

OpenStudy (anonymous):

the first one joe

OpenStudy (anonymous):

ok, so we have: \[3^{n+4} = 27^{2n}\] using the fact that: \[27 = 3^3\] we can rewrite the equation as: \[3^{n+4} =3^{3(2n)} \Rightarrow 3^{n+4} = 3^{6n} \] therefore: \[n+4 = 6n \Rightarrow 5n = 4 \Rightarrow n = \frac{4}{5}\]

OpenStudy (anonymous):

and the common base is 3...so your answer is 3^(n+4)=3^(6n) where n=4/5 by solving n+4=6n

OpenStudy (anonymous):

THANK YOU SO MUCH JOE & MATRIXX !!

OpenStudy (anonymous):

your welcome :)